OpenStudy (anonymous):

Determine whether the vectors u and v are parallel, orthogonal, or neither. u = <6, -2>, v = <8, 24>

OpenStudy (amistre64):

given a,b vectors are parallel if they are multiples of each other a , b ka,kb if they have a common ratio among parts they are parallel a/ka = 1/k b/bk = 1/k

OpenStudy (amistre64):

so what can we determine?

OpenStudy (anonymous):

It's neither

OpenStudy (amistre64):

my perp idea is flawed, parallel idea is fine tho

OpenStudy (anonymous):

Is perp = ortho?

OpenStudy (amistre64):

perp is when dot product is zero ... we dont flip nothing, they are already presented that way if perped

OpenStudy (amistre64):

yes ortho means perpendicular

OpenStudy (amistre64):

u = <6, -2> v = <8, 24> 48,-48 i was confusing flipping to perp a vector for flipping a vector

OpenStudy (anonymous):

So it is perp because they are equivalent but flipped?

OpenStudy (amistre64):

what do you recall about perpendular slopes?

OpenStudy (anonymous):

Ohhh perp slops are the opposite reciprocal

OpenStudy (amistre64):

yep flip and negate is my memory

OpenStudy (amistre64):

a slope of b/a is a vector: a,b the perp vector is some flip negate multiple of it: -kb, ka a, b -kb, ka -------- -kab, kab perp vectors have the property that the products are opposite, so their dot product is zero

OpenStudy (anonymous):

So wouldn't it be perp because 48 + -48 =0?

OpenStudy (amistre64):

yep ...